# M-LRM: Multi-view Large Reconstruction Model

### v1

@@ 4 Implementation Details and Datasets
Dataset. We use the Objaverse dataset deitke2023objaverse . We filter out 190K objects among the 800K 3D models in the dataset. We use Blender to render 32 views per object. The 3D models are scaled to fit into the cubic bounding box, whose length is 1 and center is $(0,0,0)$. We set the field-of-view to $30^{\circ}$ and the camera distance is calculated accordingly to ensure the object is well fitted at the center of the view. The first 8 views are sampled uniformly around the object with $0^{\circ}$ elevation, the second 8 views are sampled uniformly with $20^{\circ}$ elevation while the third 8 views are sampled with $-20^{\circ}$ elevation. The last 8 views are sampled randomly on the surface of the sphere. All cameras face towards $(0,0,0)$.
5 Experiments
@@
Figure 7: Visualization result of single view to 3D generation. We compare our method with Instant3D li2023instant3d , LGM tang2024lgm and TripoSR TripoSR2024 . The results of Instant3D suffer from blurry texture and severe geometry failure, while LGM tang2024lgm produces results with unreasonable color at the unseen region of the object. TripoSR can also generate visual results of high quality, but there remain many geometry artifacts. Our overall result is of better visual quality and geometry quality.
5.1 Multi-view Reconstruction
Evaluation Setting.
@@ 5.1 Multi-view Reconstruction
We compare our methods with Instant3D li2023instant3d and LGM tang2024lgm on Google Scanned Objects (GSO) downs2022google . Since the code and the model of Instant3D li2023instant3d are not available, we reproduce the model according to their paper. To fairly compare with Instant3D, we also train a tiny variant of our model that uses the same data as our reproduced Instant3D. We use the original implementation of LGM for the sparse view reconstruction task. Following existing work, we report PSNR, SSIM, and LPIPS to evaluate the visual quality of reconstruction and Chamfer distance (CD), F-Score(FS) for measuring the quality of the reconstructed mesh. We uniformly sample 16K points for measuring CD. We use two different threshold t to calculate FS for evaluation.
Results.
@@
As shown in Tab. 1, our proposed M-LRM outperforms all baselines by a large margin. Even though our Base version uses the most views compared to previous methods, our tiny version model still demonstrates the superiority of our proposed method. Compared to Instant3D, the tiny M-LRM converges much faster, and can generate images of higher fidelity. As for the LGM that uses the same number of input views, though our tiny model uses much less data than LGM, we can still generate images of higher quality, i.e. higher PSNR. However, the lack of data makes the model fail to generalize to the evaluation dataset, leading to higher LPIPS. Fig. 6 shows some visualization results of sparse-view reconstruction on the GSO dataset. Our method can not only generate novel views of better visual quality with no floaters, but also is able to extract meshes that are more plausible.
Table 1: Comparison with state-of-the-art image-to-3D counterparts on the Google Scanned Objects (GSO) dataset. $\uparrow$ means the higher the better, and $\downarrow$ indicates the lower the better.
@@
Instant3D li2023instant3d 4 256 22.18 0.958 0.142 0.0069 0.3357 0.7739
LGM tang2024lgm 4 512 21.74 0.967 0.110 0.0034 0.7108 0.9273
Ours (tiny) 4 256 24.35 0.961 0.142 0.0068 0.3357 0.771
Ours (Base) 6 512 25.23 0.977 0.096 0.0011 0.8725 0.9695
5.2 Single Image to 3D Generation
Evaluation Setting.
For single image to 3D generation, we adopt Zero123++ shi2023zero123++ to generate multiple views as conditional images. For our tiny variant and our reproduced Instant3D that takes 4 views as input, we select the first, third, fifth, and sixth images generated by Zero123++. As our model is trained with white background, while zero123++ generates multi-view images with gray background, we opt to use a refined version of zero123++ xu2024instantmesh that generates multi-view images with white background. We compare our methods with Instant3D li2023instant3d , LGM tang2024lgm , and TripoSR TripoSR2024 . For all methods, we render images at $512\times 512$ resolution.
Results.
We compare the qualitative results of our generated results with those of baseline methods. As shown in Fig. 7, we use five examples to explain our qualitative experiment. The generation outcomes of Instant3D exhibit deficiencies in texture clarity and significant faults (Cases 2 and 4) in geometric representation (cases 2, 3). Similarly, the results produced by LGM tang2024lgm display unreasonable texture color in the unobserved regions of the object (cases 1, 4, 5). Besides, the meshes produced by LGM contain some geometry artifacts. This is because of the complex pipeline for extracting mesh from 3D Gaussian representation. Although TripoSR is capable of generating visually appealing results, it still exhibits numerous imperfections in geometric accuracy (cases 1, 5) due to the lack of explicit 3D spatial modeling. In contrast, as our proposed method is geometry-aware, it yields superior visual quality and improved geometric fidelity.
5.3 Ablation Study
Table 2: The ablation study of each component.
@@ 5.3 Ablation Study
24.501 0.9746 0.1364
24.923 0.9750 0.1322
25.116 0.9765 0.1269
26.219 0.9756 0.1201
25.254 0.9772 0.1246
26.966 0.9781 0.1067
27.581 0.9803 0.0986
@@
We conduct ablation studies to demonstrate the effectiveness of our model components and architectural designs. We train these model variants with only Objaverse-LVIS subset. As shown in Tab. 2, GCA means using our proposed Multi-view Cross Attention module, GaPE means using Geometry-aware Positional Encoding, w/ Tr. Block means adding a transformer decoder block after our MCA block, and High Res. means using 64 as the resolution of the initial triplane and the geometry volume, as opposed to the original 32. When High Res. is adopted, we sample 192 points per ray in volumetric rendering instead of 96 originally.
@@
The results of the first two rows and the last two rows demonstrate the effectiveness of our proposed GCA and GaPE. As shown in the table, adding a transformer decoder block after GCA block improves the model’s performance, indicating that preserving global attention is helpful. We also show that using High Res. setting can enhance the model’s capability by a large margin.
6 Conclusion and Limitation
Conclusion.
@@
In this work, we introduce the Multi-view Large Reconstruction Model (M-LRM), which is designed to efficiently reconstruct high-quality 3D shapes from multiple views in a 3D-aware manner. Our proposed approach incorporates a multi-view consistent cross-attention scheme, enabling M-LRM to accurately extract relevant information from the input images. Additionally, we leverage the 3D priors of the input multi-view images to initialize the tri-plane tokens. By doing so, our model surpasses the capabilities of instant3D and produces 3D shapes with exceptional fidelity. Experimental results demonstrate that our model substantially improves performance and exhibits faster training convergence compared to instant3D and other LRM-series work.
Limitations and Future Work.
@@
However, we notice that there are still some limitations in our framework. (i) Currently, our model can only accept a fixed number of views as input. However, our model can be further extended to support arbitrary input views, which will be left as future work. (ii) Although our model is trained with augmented data, its performance would still inevitably be affected by inconsistent multi-view input when doing single-view to 3D generation tasks. However, we believe that this problem can be solved as the multi-view generation models advance.

### v2

@@ III-C Training Objectives
IV Implementation details
IV-A Data Preparation
Datasets. We use the Objaverse[63] dataset to train our base model, where 190K objects are filtered out among the 800K 3D models in the original dataset. We remove the objects without texture and the objects with tags indicating low quality.
For our reproduced Instant3D, tiny model and all the models in our ablation study, we opt to use Objaverse-LVIS, a subset of Objaverse, due to the limitation of computational cost. 31K objects in Objaverse-LVIS are used for training.
Commonly used dataset Google Scanned Objects(GSO)[72] is adopted as our evaluation set following previous work[65]. All 1030 objects from GSO are used for evaluation.
@@ IV-A Data Preparation
Figure 6: Visualization result of single view to 3D generation. We compare our method with Instant3D [65], LGM [59] and TripoSR [73]. The results of Instant3D suffer from blurry texture and severe geometry failure, while LGM [59] produces results with unreasonable color at the unseen region of the object. TripoSR can also generate visual results of high quality, but there remain many geometry artifacts. Our overall result is of better visual quality and geometry quality.
Rendering Strategy. We use Blender to render the images. The 3D models are scaled to fit into the cubic bounding box, whose length is 1 and center is $(0,0,0)$. We set the field-of-view to $30^{\circ}$ and the camera distance is calculated accordingly to ensure the object is well-fitted at the center of the view. For Objaverse-LVIS and GSO, we render 24 views for each object. The first 8 views are sampled uniformly around the object with $0^{\circ}$ elevation, while the second 8 views are sampled uniformly with $20^{\circ}$ elevation. The last 8 views are sampled randomly on the surface of the sphere. For our Objaverse dataset, we render 32 views for each object. In addition to the 24 views mentioned before, we render 8 extra views that are sampled uniformly around the object with elevation $-20^{\circ}$.All cameras face towards $(0,0,0)$. Uniform lighting is adopted in all render processes. All images are rendered at the resolution of $512\times 512$.
IV-B Training Scheme
It takes about 3 days to train our base M-LRM on 32 NVIDIA H800 GPU (80GB) for 20 epochs. The batch size for each GPU was set to 1, and gradient accumulation was employed for 8 steps, resulting in a total batch size of 512. At each iteration, we randomly sample 10 views among the total 32 images, where 6 of them are used as input conditional images, and the other 4 for supervision. AdamW optimizer is used with $\beta_{1},\beta_{2}$ set to $0.9,0.95$ respectively, and weight decay is set to 0.05. We increase the learning rate from 4e-10 to 4e-4 for the first epoch. After warmup, we adopt cosine annealing strategy for the learning rate and the learning rate decay drops to 0 at the end of training. The coefficients for different loss components are $\lambda_{1}=1.0$, $\lambda_{2}=0.1$ and $\lambda_{3}=1.0$. Following LGM [59], we adopt camera augmentation and image grid distortion as data augmentation. In each iteration, we randomly add noise to the input conditional camera poses(elevation, azimuth and camera distance) and apply grid distortion. The Gaussian noises are ranged in $[-10,10]$ for elevation and azimuth and $[-0.1,0.1]$ for camera distance. 50% of the iterations are affected by the data augmentation.
For other models that are trained on Objaverse-LVIS, we set the batch size to 8 for each GPU, leading to a batch size of 64 in total with 8 H800 GPUs. We train these models for 100 epochs. The other hyper-parameters are the same.
IV-C Network Architecture
To implement our base M-LRM, we utilize DINOv2-base with a patch size of 14 as our image encoder. The resolution for the 6 input images is $504\times 504$. The transformer triplane decoder consists of 6 layers with a hidden dimension of 768. We employ multi-head attention with 8 heads, and each head has a hidden dimension of 64. The learnable triplane tokens are initialized with a resolution of $64\times 64$ and 768 channels. To upsample the triplane, we use a transpose convolution layer with a kernel size of 2 and a stride of 2, resulting in a resolution of 128x128 and 40 channels. The concatenated triplane features are then decoded using a 4-layer MLP into density and RGB values. We sample 192 points per ray. The model is supervised under the resolution of $512\times 512$. At each iteration, a randomly cropped window of size $128\times 128$ is sampled for supervision for saving GPU memory.
We construct the geometry volume with grid size $V=64$, yielding a resolution of $64^{3}$. After projecting the image features into spatial volumes, we aggregate them using a 5-layer 3D convolution network with channel sizes of [64, 128, 128, 128, 64] into feature volume $\mathbf{G}$. We derive our positional embedding $\mathbf{T}_{g}$ from the volume $\mathbf{G}$ using three 2D convolution layers with a kernel size of 3 and a stride of 1.
For our reproduced Instant3D, tiny model and all the models in our ablation study, we use 4 views as input. All images are resized to $256\times 256$ for supervision. The transformer triplane decoder of our tiny model consists of 5 layers instead of 6. The decoder of our reproduced Instant3D has 10 layers of transformer blocks. Camera augmentation and grid distortion augmentation are disabled. All other settings are the same as our base model.
V Experiments
Figure 7: More visualization results of single-view generation and multi-view reconstruction. Given multiple images, our M-LRM is capable of reconstructing the object. The high fidelity novel views and high quality generated mesh demonstrates the superiority of our proposed method.Given a single image, we first synthesize the multi-view images with multi-view generation model, than feed the images into our M-LRM. Our proposed method facilitates generating novel views and meshes.
V-A Multi-view Reconstruction
@@
Evaluation Setting. We compare our methods with Instant3D [65] and LGM [59] on Google Scanned Objects (GSO) [72]. Since the code and the model of Instant3D [65] are not available, we reproduce the model according to their paper. To fairly compare with Instant3D, we also train a tiny variant of our model that uses the same data as our reproduced Instant3D. As for the comparison with LGM, we use the original implementation and default hyper-parameter setting of LGM for the sparse view reconstruction task. Following existing work, we report PSNR, SSIM, and LPIPS to evaluate the visual quality of reconstruction and Chamfer distance (CD), F-Score(FS) for measuring the quality of the reconstructed mesh. We uniformly sample 16K points for measuring CD. We use two different thresholds $t$ to calculate FS for evaluation. All images are rendered at the resolution of $512\times 512$ for evaluation.
For all 1030 objects in GSO, we render 24 views following the same setting described in the rendering strategy(Sec. 6). Four views with elevation $0^{\circ}$ and azimuth $[0,90^{\circ},180^{\circ},270^{\circ}]$ are input for all the 4-view models. For our base model, two extra views with elevation $20^{\circ}$ and azimuth $[45^{\circ},225^{\circ}]$ are used as input conditional images. The rest views are utilized to calculate the average PSNR, SSIM, and LPIPS.
@@
Results. As shown in Tab. I, our proposed M-LRM outperforms all baselines by a large margin. Even though our base version uses the most views compared to previous methods, our tiny version model still demonstrates the superiority of our proposed method. Although our tiny model is trained with much less data and lower resolution than LGM, it can still generate images and meshes of higher quality. However, the lack of data makes the model fail to generalize to the evaluation dataset, leading to higher LPIPS. When our M-LRM is equipped with more input views and more training data, it can produce 3D objects that are both visually and geometrically plausible, as demonstrated by the results of our base model.
Fig. 5 shows some visualization results of sparse-view reconstruction on the GSO dataset. Our method can not only generate novel views of better visual quality with no floaters but also is able to extract meshes that are more plausible. Instant3D and LGM generate blurry novel views and the meshes they generate suffer from severe geometric failure. This indicates that LRM’s learning ability is limited without explicit 3D priors. With our proposed method, our tiny model successfully obtains the geometric reconstruction ability with limited data. When we extend the model to more input views and more training data, our base model can reconstruct high-quality objects.
V-B Single Image to 3D Generation
TABLE II: Comparison of single-view generation with different single-view generation methods on the Google Scanned Objects(GSO) dataset.
@@ V-B Single Image to 3D Generation
Instant3D 0.0093 0.4597 0.7948
TripoSR 0.0121 0.4548 0.6879
OpenLRM 0.0060 0.5116 0.8353
LGM 0.0086 0.4650 0.7841
LGM(Zero123++) 0.0062 0.4558 0.8315
Ours(tiny) 0.0055 0.5614 0.8465
Ours(base) 0.0045 0.6148 0.8744
Evaluation Setting. For single image to 3D generation, we adopt Zero123++ [1] to generate multiple views as conditional images. For our tiny model and our reproduced Instant3D that takes 4 views as input, we select the first, third, fifth, and sixth images generated by Zero123++ following GRM[60]. We compare our methods with Instant3D [65], TripoSR [73], OpenLRM [74], and LGM [59] qualitatively(Figure. 6) and quantitatively(Table. II). For LGM that uses external multi-view generation models to obtain multi-view images for reconstruction, we additionally adopt the same Zero123++ model as ours to test LGM for fair quantitative comparison.
We adopt GSO dataset for quantitative comparison. For each object in GSO, we collect the front view with elevation $0^{\circ}$ and azimuth $0^{\circ}$ as input to generate the multi-view images and mesh. After generating multi-view images with Zero123++ and reconstructing with our M-LRM, we extract the mesh with with density field via marching cube. The generated mesh is aligned with the GT mesh by first conducting ICP algorithm and then resizing the generated mesh to the same scale as GT. We report Chamfer distance(CD) and F-Score(FS). 16K points are sampled uniformly to calculate these metrics. Two different thresholds are adopted following the evaluation setting of our multi-view reconstruction task.
Results. We compare the qualitative results of our generated results with those of baseline methods. As shown in Fig. 6, we use three examples to explain our qualitative experiment. The generation outcomes of Instant3D exhibit deficiencies in texture clarity (cases 1 and 3). OpenLRM[74] suffers from severe geometric faults(cases 1, 2, 3), because it lacks the modeling of the invisible parts of the objects. Similarly, the results produced by LGM [59] display unreasonable texture color in the unobserved regions of the object (cases 2, 3). Besides, the geometry produced by LGM contains some geometry artifacts(case 1). Although TripoSR[73] is capable of generating visually appealing results, it still exhibits numerous imperfections in geometric accuracy (case 3) and inconsistency with input image(cases 1, 2) due to the lack of explicit 3D spatial modeling. In contrast, as our proposed method is geometry-aware, it yields superior visual quality and improved geometric fidelity.
The results of quantitative comparison are shown in Table. II. Our proposed M-LRM outperforms all previous method. Since LGM also uses an external multi-view generation framework before generating the mesh, we replace the framework used by LGM with Zero123++ for fair comparison. Although it slightly improves the generated mesh of LGM due to better consistency of the generated views, our M-LRM still generates the best results. Even with limited training data, our tiny model shows fast convergence speed and strong abilities in 3D reasoning, thus generating meshes of high quality. With sufficient data, our base model demonstrates strong generation ability.
V-C Effectiveness of Model Components
TABLE III: The ablation study of each component.
@@ V-C Effectiveness of Model Components
(1) 24.501 0.9746 0.1364
(2) 24.923 0.9750 0.1322
(3) 25.116 0.9765 0.1269
(4) 25.184 0.9770 0.1233
(5) 25.254 0.9772 0.1246
(6) 26.219 0.9756 0.1201
(7) 26.588 0.9772 0.1136
(8) 26.966 0.9781 0.1067
(9) 27.581 0.9803 0.0986
We conduct ablation studies to demonstrate the effectiveness of our model components and architectural designs. We train these model variants with only Objaverse-LVIS subset. These models are evaluated on our randomly picked test set from Objaverse-LVIS. We ensure that all data in this test set do not appear in the training set. Our ablation study follows the evaluation setting of multi-view reconstruction, but uses the test split in Objaverse-LVIS. 3K objects are randomly chosen from Objaverse-LVIS as our test split.
@@
As shown in Tab. III, GCA means using our proposed Multi-view Cross Attention module, GaPE means using Geometry-aware Positional Encoding, w/ Tr. Block means adding a transformer decoder block after our MCA block, and High Res. means using 64 as the resolution of the initial triplane and the geometry volume, as opposed to the original 32. When High Res. is adopted, we sample 192 points per ray in volumetric rendering instead of 96 originally. The large in the table indicates that we extend the original Instant3D model to have the equivalent number of parameters to ours(tiny). This is to ensure fair comparison and prove that the performance gain comes from our proposed components.
@@
The results of the model (1), (2), (3) and (4) demonstrate the effectiveness of our proposed GCA and GaPE. As shown by the results of models (4) and (5), adding a transformer decoder block after GCA block improves the model’s performance, indicating that preserving global attention is helpful. We also show that using a high-resolution setting enhances the model’s capability significantly, as shown by models (1) and (6). We also prove that the performance gain of our methods does not come from the extra parameters introduced, as demonstrated by models (7) and (9). Models (6), (8), and (9) indicate that even under high-resolution setting, our proposed methods are still effective.
V-D Robustness
TABLE IV: Experiments with inaccurate camera parameters as input. We test LGM and our base model with inaccurate camera intrinsics and poses. $s$ is the pre-defined factor to control the scale of the noise added.
Method PNSR/SSIM/LPIPS(s=0) PNSR/SSIM/LPIPS(s=2) PNSR/SSIM/LPIPS(s=5) PNSR/SSIM/LPIPS(s=10)
LGM 21.74/0.9668/0.1104 21.41/0.9657/0.1137 20.34/0.9607/0.1288 18.78/0.9522/0.1458
Ours(base) 25.23/0.9773/0.0961 23.68/0.9746/0.0996 21.22/0.9672/0.1171 18.92/0.9562/0.1412
Since the GCA and GaPE modules we introduced in this paper highly rely on the condition camera intrinsic and poses, we conduct extra experiments to test our model’s robustness against inaccurate camera parameters and out-of-domain camera distributions.
Following our multi-view generation evaluation setting, we add extra random Gaussian noise to the elevation, azimuth, camera distance, and focal length of the condition cameras to test our base model and LGM. The results show that to some extent, our model is robust to the inaccurate camera information. However, we point out that the robustness of our model is still limited. Robust reconstruction with inaccurate camera parameters is still an open challenge.
As shown in Table IV, $s$ is the noise strength. For elevation and azimuth, the noise is ranged in $[-s,s]$. For camera distance, the noise is within $[-0.01s,0.01s]$. And for focal length, the noise in ranged in $[-10s,10s]$. Both LGM and our M-LRM suffer from performance drop given inaccurate camera intrinsics and poses as input. However, our method still performs better than LGM in all noise strength settings.
The results show that, to some extent, our model is robust to inaccurate camera parameters and can generalize to different camera distributions.
TABLE V: Experiment results with different camera distribution. We use cameras of different FOV and camera distance to test our base model. The results show that our model can generalize to different camera distributions although trained with fixed FOV and camera distance
. Camera Distribution PSNR SSIM LPIPS CD FS(t=0.001) FS(t=0.01) Default 25.23 0.977 0.096 0.0011 0.8725 0.9695 Random 24.86 0.976 0.101 0.0018 0.7405 0.9556
To test our base model’s generalizability when the given camera’s distribution is different from the one used for training, i.e. different FOV and camera distances, we re-render the evaluation dataset under a different setting. For every object in the GSO dataset, we choose a random FOV between 15 and 45, then select an appropriate camera distance accordingly to make sure that the object is well-fitted into the screen. The camera distance is randomly scaled with a factor ranging in $[0.9,1.2]$. Other settings are the same as our default render setting. The results show that although our base model is trained with fixed FOV and camera distance following Instant3D and LGM, our model shows strong generalizability against different camera distributions. The results are shown in Table V.
V-E Convergence Speed
Figure 8: Visualization of intermediate results during training process. Compared with existing LRM-based approaches, like Instant3D [65], our proposed GCA and GaPE remarkably reduce the convergence time and improve the reconstruction quality by a large margin. At the early stage of the training, our model shows stronger fitting ability and converges faster than the baseline method Instant3D. When trained for sufficient time, our proposed GaPE and GCA enhance the model’s reconstruction ability, yielding better generation results.
Figure 9: PSNR curve of different models on the Objaverse-LVIS validation set. Compared to the baseline method Instant3D, our proposed GCA and GaPE significantly facilitate the convergence of LRM model. The introduction of each module greatly speed up the convergence speed enhance the model’s ability when trained for the same time.
To showcase our superior convergence speed, we visualize the intermediate results during the training process of our reproduced Instant3D and our tiny model, as they are trained under the same setting. As shown in Figure 8, in both cases, our model converges significantly faster than Instant3D. At the very early stage of the training(400 iterations), our model starts to produce plausible and reasonable geometries, while Instant3D fails to generate satisfactory results until the later training stage.
Our quantitative results in Table I also showcase the fast convergence and learning ability of our M-LRM. With only Objaverse-LVIS subset, our tiny model can achieve better reconstruction results than LGM that use Objaverse dataset instead.
As the PSNR curve on validation set presented in Figure 9, our proposed GCA and GaPE significantly accelerate the convergence process. When GCA is implemented (green curve), our M-LRM demonstrates comparable generation capabilities to Instant3D (red curve) with only half the number of training iterations. Furthermore, incorporating both GCA and GaPE (blue curve) reduces the required epochs to one-third. When these models are trained for an equivalent number of epochs, our M-LRM significantly surpasses the performance of the other two models.
V-F Inference Speed
TABLE VI: Results of inference speed comparison. We compare the speed of 3D representation inference, mesh extraction and image rendering of different methods.
Method 3D Representation/s Mesh Extraction/s View Rendering/s
Instant3D 0.6 1.5 0.075
Instant3D(large) 0.8 1.5 0.075
LGM 0.2  300 0.01
Ours(tiny) 0.8 1.5 0.075
Ours(base) 1 1.5 0.08
Although our proposed methods introduced more parameters and operations, which might increase the computational cost of our model, these additional costs are not the bottleneck of the model’s inference performance. We compare the inference speed of our model and baseline methods in Table VI. Our M-LRM does inference more slowly than Instant3D due to the extra operations introduced, but the cost is still affordable. LGM’s inference speed is much faster because of the advantages of 3DGS. However, it is quite hard to extract mesh with 3DGS as representation. Although our proposed method introduces some overhead in the geometry-aware operations, we point out that this overhead is not the bottleneck of the inference performance, but brings large improvement in the generation quality. All inference speed results are tested on a single NVIDIA H800 GPU.
V-G More Visualization Results
We show more of our visualization results of both multi-view generation and single-view reconstruction(Figure 7). At the top of Figure 7, we display the 6 input views to our base model, 3 generated high-fidelity novel views and the generated mesh to visualize the results of our multi-view reconstruction. To examine the single-view generation ability of our model, we show some of the results at the bottom of Figure 7. We collect some images from the internet and use Zero123++ to generate multi-view images of the input, then feed the multi-view images to our M-LRM. The high-quality results demonstrate the superiority of our proposed method.
VI Conclusion and Limitation
@@
Conclusion. In this work, we introduce the Multi-view Large Reconstruction Model (M-LRM), which is designed to reconstruct high-quality 3D shapes from multiple views in a 3D-aware manner. Our proposed approach incorporates a multi-view consistent cross-attention scheme, enabling M-LRM to accurately extract relevant information from the input images. Additionally, we leverage the 3D priors of the input multi-view images to initialize the triplane tokens. By doing so, our model surpasses the capabilities of instant3D and produces 3D shapes with exceptional fidelity. Experimental results demonstrate that our model substantially improves performance and exhibits faster training convergence compared to instant3D and other LRM-series work.
@@
Limitations and Future Work. Although M-LRM has showcased promising performance in generating 3D objects, there are still some limitations that the current framework does not fully address. (i) Currently, our M-LRM assumes the target object is well-defined. Targets of larger scales such as rooms and open-door scenes are still challenging to reconstruct. (ii) Although our model is trained with augmented data, its performance would still inevitably be affected by inconsistent multi-view input when doing single-view to 3D generation tasks. However, we believe that this problem can be solved as the multi-view generation models advance.